OpenStudy (anonymous):

i know what the answer is but can i have an explanation. how do you solve 3y - 7 = y + 5

OpenStudy (anonymous):

@Nnesha?

Mehek (mehek14):

first subtract the y on the right with the y on the left 3y - y = ?

Mehek (mehek14):

@hieeet

OpenStudy (anonymous):

hold on sorry i was fixing the computer

OpenStudy (anonymous):

you would get 2y.

Mehek (mehek14):

ok so 2y - 7 = 5 now add 7 + 5

OpenStudy (anonymous):

2y-7=5 then add 7 and you get 12?

OpenStudy (anonymous):

then divide by 2

OpenStudy (anonymous):

and you get 6?

Mehek (mehek14):

ok so 2y = 12 now to islolate y you divide 12/2

Mehek (mehek14):

yes that's correct

OpenStudy (anonymous):

ok cool thank you!!

Mehek (mehek14):

yw